OpenStudy (anonymous):

Darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, two sisters, and him. Calculate the probability of each compound event.

OpenStudy (anonymous):

Match Term Definition (Darnell and sister), with replacement A) (mother and father), without replacement B) (sister and sister), without replacement C) (parent and sister), with replacement D)

OpenStudy (agent47):

First one with replacement is: (1/5)(1/5)=1/25

OpenStudy (kinggeorge):

There are two sister, so it should be \[\frac{1}{5}\cdot\frac{2}{5}\]

OpenStudy (agent47):

ohh yes, you're right, sorry..

OpenStudy (agent47):

2nd without replacement is: (1/5)*(1/4)=1/20

OpenStudy (anonymous):

this is one of the matching ones

OpenStudy (agent47):

Sister and sister no replacement is: (2/5)*(1/4)=1/10

OpenStudy (agent47):

Parent and sister with replacement is: (2/5)*(2/5)=4/25

OpenStudy (agent47):

Any other mistakes KG?

OpenStudy (kinggeorge):

Looks good to me.

OpenStudy (agent47):

yay!

OpenStudy (anonymous):

thanks please become a fan

OpenStudy (anonymous):

oh it didnt put the percentages so here they are for the matching A-10% B-16% C-8% & D-5%

OpenStudy (agent47):

2/25=8% 1/20=5% 1/10=10% 4/25=16%

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!!!!!!!!!